OpenStudy (anonymous):

Carl is building a model of a square pyramid for a class project. The side length of the square base is 18 inches and the slant height of the pyramid is 11 inches. What is the surface area of the model pyramid?

OpenStudy (anonymous):

511.656 in2

OpenStudy (anonymous):

\(\LARGE A=b^2 + 2b \sqrt{\frac{b^2}{4}+h^2}\), where b is the base length and h is the height. plug the values in the formula and you'll get it done :)

OpenStudy (anonymous):

lol i thought @data_lg2 is the one who's giving the answers xD